A sound wave in air has a pressure amplitude equal to 4.04 ✕ 10-3 Pa. Calculate the displacement amplitude of the wave at a frequency of 9.4 kHz. (Note: Use the following values, as needed. The equilibrium density of air is ρ = 1.20 kg/m3; the speed of sound in air is v = 343 m/s. Pressure variations ΔP are measured relative to atmospheric pressure, 1.013 105 Pa.)
Answer #1

Po = ρvAω

Where, Po = pressure amplitude, ρ = density, v = velocity, A = wave amplitude, ω = angular frequency = 2\small \pif

4.04 x 10-3 = 1.20*343*A*2\small \pi*9400

A = 1.66187 x 10-10 m